# Polarity transitions induced by symmetry-breaking outer boundary heat flux in rapidly rotating dynamos

### v1

@@ 2.3 Evaluation of spectral coefficients
for the forward-travelling fast and slow wave solutions respectively. We separate the fast and slow MAC wave parts of the general solution, which is a linear superposition of the two wave solutions (Sreenivasan & Maurya, 2021). For example,
@@ 2.4 Complementarity of vertical and horizontal buoyancies
$\lambda_{1,2}$$\approx$ ±$\pm\bigg{(}\omega_{C}+\frac{\omega_{M}^{2}}{\omega_{C}}\bigg{)}+\mathrm{i}\,\frac{\omega_{M}^{2}\omega_{\eta}}{\omega_{C}^{2}},$ (26)
$\lambda_{3,4}$$\approx$ ±$\pm\bigg{(}\frac{\omega_{M}^{2}}{\omega_{C}}+\frac{\omega_{A,V}^{2}-2\omega_{A,H}^{2}}{2\omega_{C}}\bigg{)}+\mathrm{i}\,\omega_{\eta}\,\bigg{(}1-\frac{\omega_{A,V}^{2}-\omega_{A,H}^{2}}{2\omega_{M}^{2}}\bigg{)},$ (27)
@@
following the procedure in Sreenivasan & Maurya (2021). The resultant buoyancy frequency is thus given by
@@
which imples that slow MAC waves would vanish in an unstably stratified fluid with ${\omega_{A}}^{2}<0$ as $|\omega_{A}|$ nears $|\omega_{M}|$. For $\omega_{A,V}^{2}<0$ and $\omega_{A,H}^{2}>0$, (29) indicates that the vertical and horizontal buoyancies would play a complementary role in suppressing the slow waves. While a large vertical buoyancy requires a small horizontal buoyancy to suppress these waves, a small vertical buoyancy together with a large horizontal buoyancy would produce the same result. As we shall see later in §4, the complementarity of the buoyancy frequencies places a bound on the magnitude of the horizontal buoyancy in two-component convection.
@@ 2.5 Fast and slow MAC waves under unstable stratification and a thermal wind
$Le=\dfrac{V_{M}}{2\varOmega\delta}\approx\bigg{(}\frac{\omega_{M}}{\omega_{C}}\bigg{)}_{\!0},\quad E_{\eta}=\dfrac{\eta}{2\varOmega\delta^{2}}\approx\bigg{(}\frac{\omega_{\eta}}{\omega_{C}}\bigg{)}_{\!0},$ (30a,b)
where the subscript ‘0’ represents the initial state of the perturbation. The relative intensity of horizontal buoyancy is measured by the ratio $Ra_{\ell,H}/Ra_{\ell}$, where $Ra_{\ell}$ is the resultant local Rayleigh number given by the sum of the local vertical and horizontal Rayleigh numbers,
$Ra_{\ell,V}=\dfrac{g\alpha|\beta_{y}|\delta^{2}}{2\varOmega\eta},\quad Ra_{\ell,H}=2\dfrac{g\alpha|\beta_{z}|\delta^{2}}{2\varOmega\eta}\frac{k_{y}}{k_{z}},$ (31a,b)
@@
where a truncation value of ±$\pm 5/\delta$ is used for the two wavenumbers in the computed integrals since the initial wavenumber $k_{0}=\sqrt{3}/\delta$ (Majumder et al., 2024). The evolution of blobs into columnar structures through the propagation of damped waves is evident in $y-z$ plane at $x=0$.
Figure 3: Variation of the squares of the fundamental frequencies with $Ra_{\ell,H}/Ra_{\ell}$ under constant vertical buoyancy conditions. Panels (a,c) and (b,d) correspond to cases with $-\omega_{A,V}^{2}/\omega_{M}^{2}=0.8$ and $-\omega_{A,V}^{2}/\omega_{M}^{2}=0.3$, respectively. In panels (a) and (b), $\omega_{C}^{2}$, $\omega_{M}^{2}$, and $\omega_{A}^{2}$ are plotted while panels (c) and (d) decompose $\omega_{A}^{2}$ into its vertical ($\omega_{A,V}^{2}$) and horizontal ($\omega_{A,H}^{2}$) parts. The dotted vertical lines indicate the values of $Ra_{\ell,H}/Ra_{\ell}$ where the slow MAC wave is suppressed as $|\omega_{M}|\approx|\omega_{A}|$. Panel (e) shows the suppression points of the slow wave for various strengths of $-\omega_{A,V}^{2}/\omega_{M}^{2}$. The blue and red points in panel (e) correspond to the slow wave suppression points in panels (a) and (b), respectively. (f) The axial kinetic energy $E_{k,z}$ of the fast and slow MAC waves versus $Ra_{\ell,H}/Ra_{\ell}$ where $-\omega_{A,V}^{2}/\omega_{M}^{2}=0.3$. The parameters used are $E_{\eta}=2\times 10^{-5}$, $Le=0.03$ ($\omega_{M}/\omega_{C}=0.18$), and $t/t_{\eta}=1\times 10^{-2}$.
@@
$\bar{k}_{y}=\frac{|\!|k_{y}\,\hat{u}_{z}|\!|}{|\!|\hat{u}_{z}|\!|},\quad\bar{k}_{z}=\frac{|\!|k_{z}\,\hat{u}_{z}|\!|}{|\!|\hat{u}_{z}|\!|},\quad\bar{k}=\frac{|\!|k\,\hat{u}_{z}|\!|}{|\!|\hat{u}_{z}|\!|}.$ (35a–c)
By progressively increasing $Ra_{\ell,H}$ for a fixed $Ra_{\ell,V}$ and $Le$, $\omega_{A,H}$ increases, and the resultant buoyancy frequency $\omega_{A}$ crosses the Alfvén frequency $\omega_{M}$ at certain $Ra_{\ell,H}/Ra_{\ell}$, where the slow MAC waves of frequency $\omega_{s}$ disappear. The same exercise performed with a different value of $Ra_{\ell,V}$ reveals the complementarity between $Ra_{\ell,V}/Ra_{\ell}$ and $Ra_{\ell,H}/Ra_{\ell}$ (figure 3(e)). The red and blue points in figure 3(e) indicate the value of $Ra_{\ell,H}/Ra_{\ell}$ from figure 3(a) and (b), where $\omega_{A}\approx\omega_{M}$ so that $\omega_{s}$ approaches zero. In §3, it is shown that the inequality $|\omega_{M}|>|\omega_{A}|$ represents the dipole-dominated regime while $|\omega_{M}|\approx|\omega_{A}|$ represents the transition to a multipolar state, with polarity reversals occurring in a narrow range of $Ra_{\ell,H}/Ra_{\ell}$ that lies between the dipolar and multipolar states.
@@
where the limits of the integration are set to ±$\pm 5/\delta$. While the energy of the fast MAC waves is practically unaffected by increasing horizontal buoyancy, the slow MAC wave energy goes to zero when $|\omega_{A}|$ approximately matches $|\omega_{M}|$ (figure 3(f)).
@@
The relation between the resultant local Rayleigh number $Ra_{\ell}$ and the Elsasser number $\varLambda$, given by
$\varLambda=\bigg{(}\dfrac{\omega_{M}^{2}}{\omega_{C}\omega_{\eta}}\bigg{)}_{\!0}\sim\dfrac{V_{M}^{2}}{2\varOmega\eta},$ (37)
is shown in figure 4 for the condition of vanishing slow wave axial kinetic energy for three different values of $E_{\eta}$. Each point is obtained by increasing the forcing through $Ra_{\ell}$ and noting the transition from the regime $|\omega_{M}|>|\omega_{A}|$ to the regime $|\omega_{A}|\gtrsim|\omega_{M}|$. The resultant forcing needed to suppress the slow MAC waves increases with $\varLambda$. The linear relation between $Ra_{\ell}$ and $\varLambda$ reflects the parity between $|\omega_{M}|$ and $|\omega_{A}|$ at the state of vanishing slow MAC waves. Since the relative orders of magnitude of the basic frequencies do not depend on the orientation of the gravity, rotation and magnetic field axes, the Cartesian model serves as the basis for the study of the role of equatorially anti-symmetric boundary heat flux in the nonlinear dynamo in §3, where the state of vanishing slow waves is shown to be a proxy for polarity transitions. We anticipate a self-similar behaviour for the dipole–multipole transition in the dynamo, where both $Ra_{\ell}$ and $\varLambda$ are measurable quantities.
@@ 3 Nonlinear dynamo simulations
+(×$+(\nabla\times{\bm{B}})\times{\bm{B}}+E\nabla^{2}{\bm{u}},$ (38)
@@
The modified pressure $p^{*}$ in equation (38) is given by $p+\frac{1}{2}E\,Pm^{-1}\,|\bm{u}|^{2}$. The dimensionless parameters in the above equations are the Ekman number $E=\nu/2\varOmega L^{2}$, the Prandtl number, $Pr=\nu/\kappa$, the magnetic Prandtl number, $Pm=\nu/\eta$ and the modified vertical Rayleigh number $Ra_{V}=g\alpha\beta_{s}L^{2}/2\Omega\kappa$. Here, $g$ is the gravitational acceleration, $\nu$ is the kinematic viscosity, $\kappa$ is the thermal diffusivity and $\alpha$ is the coefficient of thermal expansion. The basic state temperature profile is one of basal heating, given by $T_{0}(r)=r_{i}r_{o}/r$, where $r_{i}$ and $r_{o}$ are the inner and outer radii of the spherical shell. The velocity and magnetic fields satisfy the no-slip and electrically insulating conditions respectively at the two boundaries. The inner boundary is isothermal while a fixed heat flux at the outer boundary is given as the sum of the uniform mean and a lateral variation. The heterogeneity at the outer boundary is measured by the ratio
$q^{\star}=\frac{\text{Maximum heat flux variation}}{\text{Mean heat flux}}.$ (42)
The calculations are performed by a pseudospectral code that uses spherical harmonic expansions in the angular coordinates $(\theta,\phi)$ and finite differences in radius $r$ (Willis et al., 2007).
The parameter space in focus is that of low inertia, wherein the Rossby number based on the characteristic length scale of convection, $Ro_{\ell}$ (Christensen & Aubert, 2006) is small. Two Ekman numbers are considered, with the value of $Pm=Pr$ set to keep $Ro_{\ell}\ll 0.1$, ensuring that the simulations remain in the low-inertia regime. We explore the effect of inhomogeneous heat flux boundary conditions at the outer boundary on the dipole–multipole transition. To this end, we progressively increase the magnitude of the heat flux heterogeneity at the outer boundary for a given $Ra_{V}$. For a given $q^{\star}$, the critical vertical Rayleigh number for the onset of nonmagnetic convection, $Ra_{V,crt}$, is obtained for neutral stability of the perturbations subject to a steady mean flow and temperature field produced by the heterogeneity (Sahoo & Sreenivasan, 2017). The ratio $Ra_{V}/Ra_{V,crt}$ (table 1) is a measure of the supercriticality of convection.
The mean spherical harmonic degrees for convection and energy injection are defined by,
@@
where $E_{k}(l)$ is the kinetic energy spectrum and $E_{T}(l)$ is the spectrum obtained from the product of the transform of $u_{r}T$ and its conjugate. The total kinetic and magnetic energies in the saturated dynamo are given by the volume integrals
@@
The relative dipole field strength $f_{dip}$, which is the ratio of the mean dipole field strength to the field strength in harmonic degrees $l=$ 1–12 at the outer boundary (Christensen & Aubert, 2006), takes values $>0.35$ in all the dipole-dominated runs (see table 1).
@@ 3.1 The effect of equatorially symmetric and anti-symmetric heat flux patterns on polarity transitions
Figure 6: Evolution of dipole colatitude with magnetic diffusion time for (a) $q^{\star}=17$, $Y_{2}^{1}$ (stable dipolar), (b) $q^{\star}=18$, $Y_{2}^{1}$ (reversing), (c) $q^{\star}=20$, $Y_{2}^{1}$ (multipolar) and (d) $q^{\star}=18$ and $30$, $Y_{2}^{2}$ (stable dipolar). The other dynamo parameters are $Ra_{V}=2500$, $E=1.2\times 10^{-5}$, $Pm=Pr=1$.
@@
Figure 7: Contours of the radial magnetic field at the outer boundary for $Ra_{V}=2500$ at $q^{\star}=17$ (a), 18 (b), 20 (c); and for $Ra_{V}=25000$ at $q^{\star}=5$ (d), 6 (e), 7 (f). The other dynamo parameters are $E=1.2\times 10^{-5}$, $Pm=Pr=1$. A $Y_{2}^{1}$ heat flux heterogeneity is applied at the outer boundary.
We begin our simulations with a strong field dynamo ($B^{2}_{\text{rms}}>1$) and increase the heterogeneity of the heat flux at the outer boundary in steps while keeping all other parameters fixed. Figure 5 (a)–(d) gives examples of the equatorially anti-symmetric, symmetric and composite boundary conditions whose effects on polarity transitions are analyzed in this study. The simulations are performed for at least one magnetic diffusion time in the saturated state.
@@
For $E=1.2\times 10^{-5}$, $Pm=Pr=1$, and $Ra_{V}=2500$, a strong dipole-dominated dynamo undergoes well-defined polarity reversals as shown in figure 6(b) by progressively increasing $q^{\star}$ for the equatorially anti-symmetric $Y_{2}^{1}$ heat flux pattern. With a further increase in $q^{\star}$, the dynamo exhibits a multipolar solution (figure 6(c)). However, the equatorially symmetric $Y_{2}^{2}$ heat flux pattern does not produce polarity transitions even with a strong heterogeneity (see figure 6(d)). For the same parameters, the Earth-like composite heat flux pattern in figure 5(d), which consists of both symmetric and anti-symmetric parts, produces polarity transitions. From the evolution of the dipole axis tilt (figure 16, Appendix A), well-defined polarity reversals are noted for $q^{*}=13$. For the antisymmetric $Y_{2}^{1}$ boundary heat flux variation, the radial magnetic fields in the dipolar, reversing and multipolar states are shown in figure 7 for two series with different thermal forcing. In table 1, the runs for two values of $Ra_{V}$, $2500$ and $20000$, are presented with different heat flux patterns. In the reversing runs, the ratio of magnetic to kinetic energy falls below unity as found in earlier studies (Kutzner & Christensen, 2002; Tassin et al., 2021). The main output parameters of the polarity-reversing dynamo runs with the anti-symmetric $Y_{2}^{1}$ heat flux boundary condition are reported in table 2.
$q^{\star}$ $Ra_{V}/Ra_{V,crt}$ $N_{r}$ $l_{max}$ $Rm$ $Ro_{\ell}$ $l_{C}$ $l_{E}$ $\overline{m}$ $\bar{k}_{s}$ $\bar{k}_{z}$ $E_{k}$ $E_{m}$ $f_{dip}$ $B^{2}_{peak}$ $B^{2}_{rms}$ $Ra_{\ell,H}/Ra_{\ell}$ Type
×$\times 10^{5}$$\times 10^{5}$ D/R/M

### v2

@@ 2.3 Evaluation of spectral coefficients
for the forward-travelling fast and slow wave solutions respectively. We separate the fast and slow MAC wave parts of the general solution, which is a linear superposition of the two wave solutions (Sreenivasan and Maurya, 2021). For example,
@@ 2.4 Complementarity of vertical and horizontal buoyancies
$\lambda_{1,2}$ $\approx$ $\pm\bigg(\omega_{C}+\frac{\omega_{M}^{2}}{\omega_{C}}\bigg)+\mathrm{i}\,\frac{\omega_{M}^{2}\omega_{\eta}}{\omega_{C}^{2}},$ (28)
$\lambda_{3,4}$ $\approx$ $\pm\bigg(\frac{\omega_{M}^{2}}{\omega_{C}}+\frac{\omega_{A,V}^{2}-2\omega_{A,H}^{2}}{2\omega_{C}}\bigg)+\mathrm{i}\,\omega_{\eta}\,\bigg(1-\frac{\omega_{A,V}^{2}-\omega_{A,H}^{2}}{2\omega_{M}^{2}}\bigg),$ (29)
@@
following the procedure in Sreenivasan and Maurya (2021). Here, the slow MAC waves of frequency $\lambda_{3,4}$ are damped on the time scale $(\omega_{\eta})_{0}^{-1}=(\eta k_{0}^{2})^{-1}\sim\delta^{2}/\eta$, where the subscript ‘0’ represents the initial state of the buoyancy disturbance (see also Sreenivasan and Maurya, 2021).
From (29), the resultant buoyancy frequency is given by
@@
which implies that slow MAC waves would vanish in an unstably stratified fluid with ${\omega_{A}}^{2}<0$ as $|\omega_{A}|$ nears $|\omega_{M}|$. Since $\beta_{y}<0$ and $\beta_{z}>0$ in equation (6), $\omega_{A,V}^{2}<0$ and $\omega_{A,H}^{2}>0$ in equation (31). These two buoyancy frequencies contribute to the resultant buoyancy frequency. An increase in either $|\omega_{A,H}|$ or $|\omega_{A,V}|$ results in an increase in $|\omega_{A}|$, which may match $|\omega_{M}|$ and thereby suppress the slow MAC waves. While a large vertical buoyancy requires a small horizontal buoyancy to suppress these waves, a small vertical buoyancy together with a large horizontal buoyancy would produce the same result. As we shall see later in §4, the complementarity of the buoyancy frequencies places a bound on the magnitude of the horizontal buoyancy in two-component convection.
@@ 2.5 Fast and slow MAC waves under unstable stratification and a thermal wind
$Le=\dfrac{V_{M}}{2\varOmega\delta}\approx\bigg(\frac{\omega_{M}}{\omega_{C}}\bigg)_{\!0},\quad E_{\eta}=\dfrac{\eta}{2\varOmega\delta^{2}}\approx\bigg(\frac{\omega_{\eta}}{\omega_{C}}\bigg)_{\!0},$ (33a,b)
The relative intensity of horizontal buoyancy is measured by the ratio $Ra_{\ell,H}/Ra_{\ell}$, where $Ra_{\ell}$ is the resultant local Rayleigh number given by the sum of the local vertical and horizontal Rayleigh numbers,
$Ra_{\ell,V}=\dfrac{g\alpha|\beta_{y}|\delta^{2}}{2\varOmega\eta},\quad Ra_{\ell,H}=2\dfrac{g\alpha|\beta_{z}|\delta^{2}}{2\varOmega\eta}\frac{k_{y}}{k_{z}},$ (34a,b)
@@
where a truncation value of $\pm 5/\delta$ is used for the two wavenumbers in the computed integrals since the initial wavenumber $k_{0}=\sqrt{3}/\delta$ (Majumder et al., 2024). The evolution of blobs into columnar structures through the propagation of damped waves is evident in the yzy-z plane at $x=0$.
Figure 3: Variation of the squares of the fundamental frequencies with $Ra_{\ell,H}/Ra_{\ell}$ under constant vertical buoyancy conditions. Panels (a,c) and (b,d) correspond to cases with $\omega_{A,V}^{2}/\omega_{M}^{2}=0.8$ and $\omega_{A,V}^{2}/\omega_{M}^{2}=0.3$, respectively. In panels (a) and (b), $\omega_{C}^{2}$, $\omega_{M}^{2}$, and $\omega_{A}^{2}$ are plotted while panels (c) and (d) decompose $\omega_{A}^{2}$ into its vertical ($\omega_{A,V}^{2}$) and horizontal ($\omega_{A,H}^{2}$) parts. The dotted vertical lines indicate the values of $Ra_{\ell,H}/Ra_{\ell}$ where the slow MAC wave is suppressed as $|\omega_{M}|\approx|\omega_{A}|$. Panel (e) shows the suppression points of the slow wave for various strengths of $\omega_{A,V}^{2}/\omega_{M}^{2}$. The blue and red points in panel (e) correspond to the slow wave suppression points in panels (a) and (b), respectively. (f) The axial kinetic energy $E_{k,z}$ of the fast and slow MAC waves versus $Ra_{\ell,H}/Ra_{\ell}$ where $\omega_{A,V}^{2}/\omega_{M}^{2}=0.3$. The parameters used are $E_{\eta}=2\times 10^{-5}$, $Le=0.03$ ($\omega_{M}/\omega_{C}=0.18$), and $t/t_{\eta}=1\times 10^{-2}$.
@@
[…]
By progressively increasing $Ra_{\ell,H}$ for a fixed $Ra_{\ell,V}$ and $Le$, $\omega_{A,H}$ increases, and the resultant buoyancy frequency $\omega_{A}$ crosses the Alfvén frequency $\omega_{M}$ at certain $Ra_{\ell,H}/Ra_{\ell}$, where the slow MAC waves of frequency $\omega_{s}$ disappear ($\omega_{s}=\omega_{3,4}$; see (29)). The same exercise performed with a different value of $Ra_{\ell,V}$ reveals the complementarity between $Ra_{\ell,V}/Ra_{\ell}$ and $Ra_{\ell,H}/Ra_{\ell}$ (figure 3(e)). The red and blue points in figure 3(e) indicate the value of $Ra_{\ell,H}/Ra_{\ell}$ from figure 3(a) and (b), where $|\omega_{A}|\approx|\omega_{M}|$ so that $\omega_{s}$ approaches zero. In §3, it is shown that the inequality $|\omega_{M}|>|\omega_{A}|$ represents the dipole-dominated regime while $|\omega_{M}|\approx|\omega_{A}|$ represents the transition to a multipolar state, with polarity reversals occurring in a narrow range of $Ra_{\ell,H}/Ra_{\ell}$ that lies between the dipolar and multipolar states.
@@
where the limits of the integration are set to $\pm 5/\delta$. While the energy of the fast MAC waves is practically unaffected by increasing horizontal buoyancy, the slow MAC wave energy goes to zero when $|\omega_{A}|$ approximately matches $|\omega_{M}|$ (figure 3(f)).
@@
The relation between the resultant local Rayleigh number $Ra_{\ell}$ and the ratio of frequencies given by
$\bigg(\dfrac{\omega_{M}^{2}}{\omega_{C}\omega_{\eta}}\bigg)_{\!0}\sim\dfrac{V_{M}^{2}}{2\varOmega\eta}$ (40)
is shown in figure 4 for the condition of vanishing slow wave axial kinetic energy for three different values of $E_{\eta}$. Each point is obtained by increasing the forcing through $Ra_{\ell}$ and noting the transition from the regime $|\omega_{M}|>|\omega_{A}|$ to the regime $|\omega_{A}|\gtrsim|\omega_{M}|$. The resultant forcing needed to suppress the slow MAC waves increases with $(\omega_{M}^{2}/\omega_{C}\omega_{\eta})_{0}$, which is a measure of the square of the peak dimensionless magnetic field in the dynamo (see §3.4 below). The linear relation obtained in figure 4 reflects the parity between $|\omega_{M}|$ and $|\omega_{A}|$ at the state of vanishing slow MAC waves. Since the relative orders of magnitude of the basic frequencies do not depend on the orientation of the gravity, rotation and magnetic field axes, the Cartesian model serves as the basis for the study of the role of equatorially anti-symmetric boundary heat flux in the nonlinear dynamo in §3, where the state of vanishing slow waves is shown to be a proxy for polarity transitions. We anticipate a self-similar behaviour for the dipole–multipole transition in the dynamo, where both $Ra_{\ell}$ and the square of the peak magnetic field are measurable quantities.
@@ 3 Nonlinear dynamo simulations
$+(\nabla\times{\bm{B}})\times{\bm{B}}+E\nabla^{2}{\bm{u}},$ (41)
@@
The modified pressure $p^{*}$ in equation (41) is given by $p+\frac{1}{2}E\,Pm^{-1}\,|\bm{u}|^{2}$. The dimensionless parameters in the above equations are the Ekman number $E=\nu/2\varOmega L^{2}$, the Prandtl number, $Pr=\nu/\kappa$, the magnetic Prandtl number, $Pm=\nu/\eta$ and the modified vertical Rayleigh number $Ra_{V}=g\alpha\beta_{s}L^{2}/2\Omega\kappa$. Here, $g$ is the gravitational acceleration, $\nu$ is the kinematic viscosity, $\kappa$ is the thermal diffusivity and $\alpha$ is the coefficient of thermal expansion. The basic state conductive temperature profile is one of basal heating, given by $T_{0}(r)=r_{i}r_{o}/r$, where $r_{i}$ and $r_{o}$ are the inner and outer radii of the spherical shell. The velocity and magnetic fields satisfy the no-slip and electrically insulating conditions respectively at the two boundaries. The inner boundary is isothermal while a fixed heat flux at the outer boundary is given as the sum of the uniform mean and a lateral variation. The heterogeneity at the outer boundary is measured by the ratio
$q^{*}=\frac{\text{Maximum heat flux variation}}{\text{Mean heat flux}}.$ (45)
The mean basic state heat flux in the denominator of (45) represents the superadiabatic heat flux from the core (e.g. Olson, 2015, p. 11). The maximum heat flux variation at the outer boundary represents the magnitude of the peak-to-peak variation at the CMB relative to the mean superadiabatic value. The calculations are performed by a pseudospectral code that uses spherical harmonic expansions in the angular coordinates $(\theta,\phi)$ and finite differences in radius $r$ (Willis et al., 2007).
The parameter space in focus is that of low inertia, wherein the Rossby number based on the characteristic length scale of convection, $Ro_{\ell}$ (Christensen and Aubert, 2006) is small. Two Ekman numbers are considered, with the value of $Pm=Pr$ set to keep $Ro_{\ell}\ll 0.1$, ensuring that the simulations remain in the low-inertia regime. We explore the effect of inhomogeneous heat flux boundary conditions at the outer boundary on the dipole–multipole transition. To this end, we progressively increase the magnitude of the heat flux heterogeneity at the outer boundary for a given $Ra_{V}$. For a given $q^{*}$, the critical vertical Rayleigh number for the onset of nonmagnetic convection, $Ra_{V,\,c}$, is obtained for neutral stability of the perturbations subject to a steady mean flow and temperature field produced by the heterogeneity (Sahoo and Sreenivasan, 2017). The ratio $Ra_{V}/Ra_{V,\,c}$ (table 1) is a measure of the supercriticality of convection.
The mean spherical harmonic degrees for convection and energy injection, denoted by $l_{c}$ (Christensen and Aubert, 2006) and $l_{E}$ (Sreenivasan et al., 2014; Varma and Sreenivasan, 2022; Majumder et al., 2024) respectively, are given by,
@@
where $E_{k}(l)$ is the kinetic energy spectrum and $E_{T}(l)$ is the spectrum obtained from the product of the transform of $u_{r}T$ and its conjugate, showing the spectral distribution of scales at which energy is injected by buoyancy. The total kinetic and magnetic energies in the saturated dynamo are given by the volume integrals
@@
The relative dipole field strength $f_{dip}$, which is the ratio of the mean dipole field strength to the field strength in harmonic degrees $l=$ 1–12 at the outer boundary (Christensen and Aubert, 2006), takes values >0.35>0.35 in all the dipole-dominated runs (see table 1).
The square of the peak magnetic field, denoted by $B^{2}_{peak}$ (tables 1 & 2), is obtained from the time-averaged value of the magnetic field at the peak-field location in the saturated state of each run.
@@ 3.1 The effect of equatorially symmetric and anti-symmetric heat flux patterns on polarity transitions
Figure 6: Evolution of dipole colatitude with magnetic diffusion time for (a) $q^{*}=17$, $Y_{2}^{1}$ (stable dipolar), (b) $q^{*}=18$, $Y_{2}^{1}$ (reversing), (c) $q^{*}=20$, $Y_{2}^{1}$ (multipolar) and (d) $q^{*}=18$ and $30$, $Y_{2}^{2}$ (stable dipolar). The other dynamo parameters are $Ra_{V}=2500$, $E=1.2\times 10^{-5}$, $Pm=Pr=1$.
@@
Figure 7: Contours of the radial magnetic field at the outer boundary for $Ra_{V}=2500$ at $q^{*}=17$ (a), 18 (b), 20 (c); and for $Ra_{V}=25000$ at $q^{*}=5$ (d), 6 (e), 7 (f). The other dynamo parameters are $E=1.2\times 10^{-5}$, $Pm=Pr=1$. A $Y_{2}^{1}$ heat flux heterogeneity is applied at the outer boundary.
We begin our simulations with a strong field dynamo ($B^{2}_{rms}>1$) and increase the heterogeneity of the heat flux at the outer boundary in steps while keeping all other parameters fixed. Figure 5 (a)–(d) gives examples of the equatorially anti-symmetric, symmetric and composite boundary conditions whose effects on polarity transitions are analyzed in this study. The simulations are performed for at least one magnetic diffusion time in the saturated state.
@@
For $E=1.2\times 10^{-5}$, $Pm=Pr=1$, and $Ra_{V}=2500$, a strong dipole-dominated dynamo undergoes well-defined polarity reversals as shown in figure 6(b) by progressively increasing $q^{*}$ for the equatorially anti-symmetric $Y_{2}^{1}$ heat flux pattern. With a further increase in $q^{*}$, the dynamo exhibits a multipolar solution (figure 6(c)). However, the equatorially symmetric $Y_{2}^{2}$ heat flux pattern does not produce polarity transitions even with a strong heterogeneity (see figure 6(d)). For the same parameters, the Earth-like composite heat flux pattern in figure 5(d), which consists of both symmetric and anti-symmetric parts, produces polarity transitions. From the evolution of the dipole axis tilt (figure 17, Appendix A), well-defined polarity reversals are noted for $q^{*}=13$. For the anti-symmetric $Y_{2}^{1}$ boundary heat flux variation, the radial magnetic fields in the dipolar, reversing and multipolar states are shown in figure 7 for two series with different vertical thermal forcing. In table 1, the runs for two values of $Ra_{V}$, $2500$ and $20000$, are presented with different heat flux patterns. In the reversing runs, the ratio of magnetic to kinetic energy $E_{m}/E_{k}$ falls below unity, as noted in earlier studies (Kutzner and Christensen, 2002; Tassin et al., 2021) although a recent study (Frasson et al., 2025) suggests that multipolar solutions can exist for $E_{m}/E_{k}>1$. Reversing dynamo models in general have not realized $E_{m}/E_{k}>1$ probably because their magnetic Ekman number $E_{\eta}$ is much higher than that in the Earth’s core in the relatively small length scales. In contrast, the criterion of vanishing slow MAC wave helicity considered in this study applies to the energy-containing scales where density perturbations are continually produced, and is independent of the choice of $E_{\eta}$ in the low-inertia limit that rapidly rotating planets operate in.
The main output parameters of the polarity-reversing dynamo runs with the anti-symmetric $Y_{2}^{1}$ heat flux boundary condition are reported in table 2.
$q^{*}$ $\dfrac{Ra_{V}}{Ra_{V,\,c}}$ $N_{r}$ $l_{max}$ RmRm $Ro_{\ell}$ $l_{C}$ $l_{E}$ $\overline{m}$ $\bar{k}_{s}$ $\bar{k}_{z}$ $E_{k}$ $E_{m}$ $f_{dip}$ $B^{2}_{peak}$ $B^{2}_{rms}$ $\bigg|\dfrac{\omega_{A}}{\omega_{M}}\bigg|$ $\dfrac{Ra_{\ell,H}}{Ra_{\ell}}$ Type
$\times 10^{5}$ $\times 10^{5}$ D/R/M